# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ III.2 Fermionic self-energy and momentum distribution
multiplied by the number of composite fermions $n_{\mathrm{CF}}$. The second term in Eq. (41) is clearly associated with the composite fermion contribution to $n_{\mathrm{F}}({\mathbf{k}})$ whereas the first is due to dressed unpaired fermions as evidenced by the presence of a Fermi step function with argument $-\tilde{\xi}_{\mathbf{k}}^{\rm F}$ and a quasi-particle weight in front of it.
We now deal with the second term in Eq. (III.2). It has a pole in $z=\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}$ which contributes
$n_{\mathrm{F}}^{(2)}({\mathbf{k}})=\frac{\Delta_{0}^{2}}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}})^{2}}\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}),$ (43)
and a double pole in $z=\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}$ which contributes
$n_{\mathrm{F}}^{(3)}({\mathbf{k}})=-\frac{\Delta_{0}^{2}}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}})^{2}}\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}).$ (44)
These two terms stem from the second diagram in Fig. 2(b) which favors both pairing with the condensate and atom-molecule hybridization via the condensate. On physical grounds, for ${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$ one expects to find weakly interacting atomic fermions, whereas in the opposite limit of ${n_{\mathrm{B}}}\approx{n_{\mathrm{F}}}$ one should find the specular situation of weakly interacting molecules. In these two opposite limits one has indeed $\tilde{\mu}_{\rm F}>0$, $\tilde{\mu}_{\rm CF}<0$ for $x\to 0$ and $\tilde{\mu}_{\rm CF}>0$, $\mu_{\rm F}<0$ for $x\to 1$. Therefore, expression (43) is non-zero for $x\to 1$, with $\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})=1$, whereas expression (44) is non-zero for $x\to 0$, with $\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})=1$, and is to be summed to the same type of contribution in expression (41).
To sum up, the full fermionic momentum distribution ${n_{\mathrm{F}}}({\mathbf{k}})$ is split as the sum of a contribution ${n_{\mathrm{CF}}}|\phi({\mathbf{k}})|^{2}$ from fermions bound into molecules, and an unpaired fermion contribution $n_{\mathrm{UF}}({\mathbf{k}})$ as follows:
$n_{\mathrm{F}}({\mathbf{k}})={n_{\mathrm{CF}}}|\phi({\mathbf{k}})|^{2}+n_{\mathrm{UF}}({\mathbf{k}}),\hskip 73.97733pt$ (45)
where
@@
$n_{\mathrm{UF}}({\mathbf{k}})=$ $\left[1-{n_{\mathrm{CF}}}|\phi({\mathbf{k}})|^{2}-\frac{\Delta_{0}^{2}}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})^{2}}\right]\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}),\quad$ $x\to 0$ (46a)
$n_{\mathrm{UF}}({\mathbf{k}})=$ $\frac{\Delta_{0}^{2}}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})^{2}}\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}),\quad$ $x\to 1.$ (46b)
Expressions (46a) and (46b) describe unpaired fermions in two different regimes. In Eq. (46a) atoms are the majority species and most of the fermions are unpaired because of the small number of bosons to pair with. When one inserts the contribution (46a) in Eq. (45) one recovers a nearly filled Fermi sphere, with a small depletion, described by the last term in Eq. (46a). In Eq. (46b) molecules are instead the majority species; the few unpaired fermions are strongly renormalized by interaction with molecules (since $\Delta_{0}/E_{\rm F}\ll 1$ and the denominator in Eq. (46b) never vanishes) and partially occupy single-particle states within a Fermi sphere determined by $\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})$.
Finally, we consider evaluating the frequency integration of the Green’s function (32) to obtain $n_{\mathrm{F}}({\mathbf{k}})$ in a closed form without expanding in $\Delta_{0}/{E_{\mathrm{F}}}$. This can be done analytically on the complex plane by first solving the cubic equation for the three poles $z_{i},\;i=1,2,3$
$z_{i}-\xi_{\mathbf{k}}^{\mathrm{F}}-\frac{\Delta^{2}_{0}}{z_{i}-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}-{\Delta^{2}_{\mathrm{CF}}\over z_{i}+\xi^{\mathrm{B}}_{{\mathbf{k}}}}=0,$ (47)
and then applying the residue theorem on a contour running along the imaginary axis and closed on the left-hand half of the plane, yielding
The total number density $n_{\mathrm{F}}$ is then evaluated numerically via Eq. (16).
III.3 Hugenholtz-Pines condition for condensed bosons
@@
In the absence of boson-boson repulsion ($\Sigma_{11}=\Sigma_{12}=0$) and in the strong coupling limit of interest here, the bosonic self-energy (10) acquires the form
@@
$\Sigma_{\mathrm{B}}({\mathbf{k}},\omega)=\left(2\pi{\varepsilon_{0}}\over m_{r}\right)$ $\int\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}{\Theta(-E_{\mathbf{P}}^{+})-\Theta(-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}})\over E_{\mathbf{P}}^{+}-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}}-i\omega}\right.$
$+v_{\mathbf{P}}^{2}\left.{\Theta(-E_{\mathbf{P}}^{-})-\Theta(-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}})\over E_{\mathbf{P}}^{-}-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}}-i\omega}\right],$ (49)
after a contour integration on the left-hand side of the complex frequency plane.
@@
This expression is not analytically tractable except when it is evaluated at zero momentum and frequency in the Hugenholtz-Pines condition (18), as shown in Appendix D.
@@
III.4 Complete system of semi-analytical equations in closed form and their solutions
@@
Collecting the analytic results obtained so far, it turns out that the system of Eqs. (16-18) becomes analytically tractable in the strong coupling limit. Equations (16 and 18) are given in closed form in Eq. (47) and in Eq. (88) in Appendix D respectively, while Eq. (17) does not have an analytic form. Nevertheless, for large $g$ all bosons out of the condensate are expected to bind to fermions (as $n_{\mathrm{B}}\leq n_{\mathrm{F}}$) hence allowing one to substitute Eq. (17) for bosons with Eq. (III.1) for composite fermions since $n_{\mathrm{B}}-n_{0}\simeq n_{\mathrm{CF}}$. This replacement is not just a mathematical rearrangement but reflects the physical fact that the system is now effectively a Fermi-Fermi mixture of molecules and unpaired fermions. As a consequence, one obtains the following set of equations for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$ (and implicitly $\mu_{\mathrm{CF}}\equiv\mu_{\mathrm{F}}+\mu_{\mathrm{B}}+{\varepsilon_{0}}$):
@@
  • Composite-fermion number Eq. (III.1)
    $n_{\mathrm{B}}-n_{0}=\int\frac{d^{2}P}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}\Theta(-E_{+})+v_{\mathbf{P}}^{2}\Theta(-E_{-})\right],$ (51)
    whose analytical form is given in Eq. 85;
@@
    ${\mu_{\mathrm{B}}}=\Sigma_{\mathrm{B}}({\mathbf{k}}=\mathbf{0},\omega=0)=\left(2\pi{\varepsilon_{0}}\over m_{r}\right)\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}$
    ×$\times\left[{u_{\mathbf{P}}^{2}\over E_{\mathbf{P}}^{+}-\xi^{\mathrm{F}}_{{\mathbf{P}}}}\Theta(-E_{\mathbf{P}}^{+})+{v_{\mathbf{P}}^{2}\over E_{\mathbf{P}}^{-}-\xi^{\mathrm{F}}_{{\mathbf{P}}}}\Theta(-E_{\mathbf{P}}^{-})\right],$ (52)
    whose analytic expression is given in Eq. 88.
@@
These equations can be solved with a standard root finder, and no particular care has to be taken when performing the related momentum integrations.
@@
Figure 3: (a): Chemical potentials ${\mu_{\mathrm{F}}}$, ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ and ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma_{\mathrm{CF}}$ as functions of concentration $x$, for strong BF attraction $g=2.8$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$. (b): Square of the hybridization energy $\Delta_{0}$ (in units of $E_{\rm F}^{2}$) as a function of the boson concentration $x$, for strong BF attraction $g=2.8$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$.
We now discuss their solutions in the form of the chemical potentials ${\mu_{\mathrm{F}}}$ and ${\mu_{\mathrm{B}}}$, the hybridization energy $\Delta_{0}$ and the Fermi momentum distribution profile $n_{\mathrm{F}}({\mathbf{k}})$.
In Fig. 3(a) the fermion chemical potential ${\mu_{\mathrm{F}}}$ and the subleading correction to the boson chemical potential ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ are shown along with the (shifted) composite fermion chemical potential ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma_{\mathrm{CF}}$. Trivially, the bosonic chemical potential ${\mu_{\mathrm{B}}}$ follows the dominant energy scale of the binding energy $-{\varepsilon_{0}}$, implying that as soon as a boson is added to the system it binds to a fermion (${n_{\mathrm{B}}}\leq{n_{\mathrm{F}}}$). Depending on the sign of the chemical potentials, we distinguish two clearly different yet specular regimes separated by a crossover region roughly given by the range $0.7\lesssim x\lesssim 0.9$ (shaded region). In line with what obtained in Sec. III.2, we confirm that at the left of this region (${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$), the system is made of weakly interacting atomic fermions (${\mu_{\mathrm{F}}}>0$) and strongly renormalized molecules (${\tilde{\mu}_{\mathrm{CF}}}<0$) whereas at the right of this region (${n_{\mathrm{B}}}\simeq{n_{\mathrm{F}}}$), in a specular manner, degenerate molecules (${\tilde{\mu}_{\mathrm{CF}}}>0$) and strongly dressed atomic fermions (${\mu_{\mathrm{F}}}<0$) coexist.
In the crossover region, all chemical potentials are comparable in size and on the verge of changing sign: here atoms and molecules are in a state of maximal hybridization as reflected in Fig. 3(b) where the hybridization energy squared $\Delta_{0}^{2}/{E_{\mathrm{F}}}^{2}=2{\varepsilon_{0}}n_{0}/{E_{\mathrm{F}}}{n_{\mathrm{F}}}$ defined in Sec. III.1 is reported. The inversion of behavior of $\Delta_{0}$ as a function of $x$ approximately matches the inversion of sign of the chemical potentials, being peaked exactly when ${\mu_{\mathrm{F}}}\simeq 0$. In the inset of Fig. 3(b) the condensate density $n_{0}$ is also displayed to show its exponentially suppressed value, compensated by the exponential increase of the binding energy in the resulting $\Delta_{0}$.
@@
Figure 4: Dispersions $E_{\mathbf{P}}^{\pm}$ of the poles of the strong-coupling limit $T$-matrix $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$ and bare molecular and atomic Fermi dispersions $\xi_{\mathbf{P}}^{\rm CF}$ and $\xi_{\mathbf{P}}^{\rm F}$, respectively. All data are obtained for strong BF attraction $g=2.8$ and several bosonic concentrations: (a) $x=0.95$, (b) $x=0.76$, (c) $x=0.5$, (d) $x=0.175$.
@@
Let us now analyze in detail the effect of the hybridization scale $\Delta_{0}$ through the energy spectrum $E_{\mathbf{P}}^{\pm}$ of the molecular quasi-particles defined by Eq. (25) of Sec. III.1. In Fig. 4 the dispersions $E_{\mathbf{P}}^{\pm}$ are compared with those of the bare (molecular and atomic) fermions for coupling $g=2.8$ and different values of the boson concentration $x$. At exactly matched densities ($x=1$) all fermions are essentially paired with all bosons ($n_{0}\approx 0$) and the system is effectively made of a gas of non-interacting molecules, with the bare dispersions coinciding with the dressed one since $\Delta_{0}\simeq 0$.
@@
As soon as a population imbalance sets in $x$ (see the case $x=0.95$ in Fig. 4(a)), the hybridization energy scale $\Delta_{0}$ sharply rises and hybrid quasi-particles form out of the molecular and unpaired Fermi states. This phase approximately corresponds to the extreme right of Fig. 3(a)-(b). As population imbalance further increases, even though $n_{0}$ remains exponentially suppressed ($n_{0}\propto\Delta_{0}^{2}/{\varepsilon_{0}}$), the presence of a non-zero $\Delta_{0}$ stabilizes the system by lowering the energy dispersion $E_{\mathbf{P}}^{-}$ with respect to the bare one $\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}$. This effect is rooted in the kinetic energy cost when filling up the paired and unpaired Fermi spheres: in order to reduce it, the molecular and atomic wave-functions hybridize quantum-mechanically by overlapping with the condensate wave-function (thus gaining in delocalization energy). The bottom panels of Fig. 4 display the typical level crossing (dotted lines) and avoided crossing (dashed lines) of a non-hybridized and hybridized two-level system, respectively. At the crossover concentration $x\simeq 0.7$, the hybridization is about maximal and the character of the occupied states (bottom pole dispersion) switches from molecular to mostly atomic as $x$ is reduced (with an effective mass approaching $m_{\mathrm{F}}$). At small concentration ($x=0.175$) the dispersion of occupied states approaches the limit of the weakly interacting Fermi gas $\xi_{\mathbf{k}}^{\mathrm{F}}$ (dressed impurity limit) with vanishing quasi-particle weight $v_{\mathbf{P}}^{2}\simeq 0$, in line with the extreme left region of Fig. 3(a)-(b).
@@
Figure 5: Quasi-particle weight $v_{\mathbf{P}}^{2}$ of Eq. (27) corresponding to the momentum distribution of the hybridized composite fermions for BF coupling $g=2.8$ and concentrations matching those in Fig. 4, along with the case $x=1$.

### v2

@@ 3.2 Fermionic self-energy and momentum distribution
is the bound-state wave-function for the two-body problem in vacuum, while
@@
$n_{\mathrm{UF}}({\mathbf{k}})=\left[1-{n_{\mathrm{CF}}}|\phi({\mathbf{k}})|^{2}-\frac{\Delta_{0}^{2}\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})^{2}}\right]\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})+\frac{\Delta_{0}^{2}\Theta(-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}})}{(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}})^{2}}\Theta(\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}).$ (43)
with $\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}\equiv\xi_{\mathbf{k}}^{\mathrm{F}}+2\pi n_{\rm CF}/m_{r}$. The first term in Eq. (41) corresponds to fermions bound in molecules, while the second one corresponds to unpaired fermions, as discussed in Appendix D.2.
3.3 Hugenholtz-Pines condition for condensed bosons
@@
In the absence of boson-boson repulsion ($\Sigma_{11}=\Sigma_{12}=0$) and in the strong-coupling limit of interest here, the bosonic self-energy (11) acquires the form
@@
$\Sigma_{\mathrm{B}}({\mathbf{k}},\omega)=\left(2\pi{\varepsilon_{0}}\over m_{r}\right)\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}{\Theta(-E_{\mathbf{P}}^{+})-\Theta(-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}})\over E_{\mathbf{P}}^{+}-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}}-i\omega}+v_{\mathbf{P}}^{2}{\Theta(-E_{\mathbf{P}}^{-})-\Theta(-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}})\over E_{\mathbf{P}}^{-}-\xi^{\mathrm{F}}_{{\mathbf{P}}-{\mathbf{k}}}-i\omega}\right],$ (44)
when $T(\textbf{P},\Omega)$ is replaced by $T_{\mathrm{SC}}(\textbf{P},\Omega)$ in Eq. (9) and the frequency integral is performed.
@@
The above integral can be solved analytically when, as required in the Hugenholtz-Pines condition (19), it is evaluated at zero momentum and frequency, as shown in Appendix D.
@@
3.4 Complete system of semi-analytic equations in closed form and their solutions
@@
Collecting the analytic results obtained so far, it turns out that the system of Eqs. (17)-(19) becomes analytically tractable in the strong-coupling limit. Equations (17) and (19) are given in closed form in Eq. (39) and in Eq. (D.4) of Appendix D, respectively. Equation (18) can be dealt with by observing that for large $g$ all non-condensed bosons are expected to bind with fermions into molecules (we recall that $n_{\mathrm{B}}\leq n_{\mathrm{F}}$). One can then substitute Eq. (18) for bosons with Eq. (33) for composite fermions since $n_{\mathrm{B}}-n_{0}\simeq n_{\mathrm{CF}}$. As a consequence, one obtains the following set of equations for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$ (and implicitly $\mu_{\mathrm{CF}}\equiv\mu_{\mathrm{F}}+\mu_{\mathrm{B}}+{\varepsilon_{0}}$):
@@
  • Composite-fermion number Eq. (33)
    $n_{\mathrm{B}}-n_{0}=\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}\Theta(-E_{+})+v_{\mathbf{P}}^{2}\Theta(-E_{-})\right],$ (46)
    whose analytic form is given in Eq. (D.1);
@@
    ${\mu_{\mathrm{B}}}=\Sigma_{\mathrm{B}}(\mathbf{0},0)=\left(2\pi{\varepsilon_{0}}\over m_{r}\right)\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[{u_{\mathbf{P}}^{2}\over E_{\mathbf{P}}^{+}-\xi^{\mathrm{F}}_{{\mathbf{P}}}}\Theta(-E_{\mathbf{P}}^{+})+{v_{\mathbf{P}}^{2}\over E_{\mathbf{P}}^{-}-\xi^{\mathrm{F}}_{{\mathbf{P}}}}\Theta(-E_{\mathbf{P}}^{-})\right],$ (47)
    whose analytic expression is given in Eq. (D.4).
@@
This simple system of equations can be solved for $\mu_{\rm F},\mu_{\rm B}$, and $n_{0}$ with a standard root finder, and no particular care has to be taken when performing the related momentum integrals.
@@
Figure 3: (a): Chemical potentials ${\mu_{\mathrm{F}}}$, ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ and ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma^{0}_{\mathrm{CF}}$ (in units of $E_{\rm F}$) as functions of concentration $x$, for strong BF attraction $g=2.8$ and BB repulsion $\eta_{\mathrm{B}}=0$. (b): Square of the hybridization energy $\Delta_{0}^{2}=2\pi{\varepsilon_{0}}n_{0}/m_{r}$ (in units of $E_{\rm F}^{2}$) as a function of the boson concentration $x$, for strong BF attraction $g=2.8$ and BB repulsion $\eta_{\mathrm{B}}=0$. Insets: Condensate density $n_{0}$ in units of $n_{\rm F}$ (top) and condensate fraction $x$ (bottom) for the same set of parameters.
Figures 3--6 present the results for different physical quantities based on the solutions of Eqs. (45)-(47). In all calculations, we will consider equal boson and fermion masses: $m_{\rm B}=m_{\rm F}$.
In Fig. 3(a) the fermion chemical potential ${\mu_{\mathrm{F}}}$ and the boson chemical potential ${\mu_{\mathrm{B}}}$ (once subtracted the leading term $-{\varepsilon_{0}}$) are shown along with the shifted composite fermion chemical potential ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma^{0}_{\mathrm{CF}}$. Trivially, the bosonic chemical potential ${\mu_{\mathrm{B}}}$ follows the dominant energy scale of the binding energy $-{\varepsilon_{0}}$, implying that as soon as a boson is added to the system it binds with a fermion (recalling that ${n_{\mathrm{B}}}\leq{n_{\mathrm{F}}}$).
Two clearly different regimes can be distinguished in the opposite limits $x\to 0$ and $x\to 1$. For $x\simeq 0$, the system is made of weakly interacting atomic fermions and composite fermions (molecules) which are strongly affected by interaction. One sees indeed that ${\mu_{\mathrm{F}}}$ approaches the non-interacting value $E_{\rm F}$. At the same time, ${\tilde{\mu}_{\mathrm{CF}}}$ is negative, implying that $\Theta(-\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}})=0$ and a non-vanishing composite fermion density is obtained only because the composite fermion momentum distribution ${n_{\mathrm{CF}}}({\mathbf{P}})=v_{{\mathbf{P}}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ strongly differs from the ideal Fermi distribution $\Theta(-\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}})$ (see also Fig. 5 below), indicating strong interaction effects on the molecules. The reverse situation occurs for $x\to 1$. In this case ${\mu_{\mathrm{F}}}<0$, implying that $\Theta(-\xi_{\mathbf{k}}^{\mathrm{F}})=0$, while ${\tilde{\mu}_{\mathrm{CF}}}$ approaches $E_{\rm F}/2$, the expected value for a filled Fermi sphere of non-interacting molecules of mass $M=2m_{\rm F}$ (note also that $\Sigma^{0}_{\rm CF}=0$ in this case, implying ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}$).
We therefore identify a crossover region $0.7\lesssim x\lesssim 0.9$ (shaded area in Fig. 3) in which the two chemical potentials ${\mu_{\mathrm{F}}}$ and ${\tilde{\mu}_{\mathrm{CF}}}$ cross and change their sign (somewhat analogously with the use of the chemical potential $\mu_{\rm F}$ as a proxy for the crossover region when studying the BCS-BEC crossover [111, 112]). In this region, atoms and molecules are in a state of maximal hybridization as reflected in Fig. 3(b) reporting the square of the hybridization energy $\Delta_{0}^{2}/{E_{\mathrm{F}}}^{2}=2\pi{\varepsilon_{0}}n_{0}/m_{r}$ introduced in Sec. 3.1.
@@
Figure 4: Hybridized dispersions $E_{\mathbf{P}}^{\pm}$ of the poles of the strong-coupling limit $T$-matrix $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$ and unhybridized molecular and atomic dispersions $\tilde{\xi}_{\mathbf{P}}^{\rm CF}$ and $\xi_{\mathbf{P}}^{\rm F}$ (in units of $E_{\rm F}$) as function of $P/k_{\mathrm{F}}$. All data are obtained for strong BF attraction $g=2.8$ and BB repulsion $\eta_{\mathrm{B}}=0$. Different bosonic concentrations are considered: (a) $x=0.95$, (b) $x=0.76$, (c) $x=0.5$, (d) $x=0.175$.
In the inset of Fig. 3(b) (upper panel) the condensate density $n_{0}$ is also displayed, to show that its value is exponentially suppressed for strong coupling. This exponentially small value is however compensated by the exponentially large binding energy in the expression for $\Delta_{0}$, yielding the sizable values of $\Delta_{0}$ shown in the main panel of Fig. 3(b). For completeness, the inset of Fig. 3(b) (lower panel) also reports the condensate fraction $x$ as a function of concentration for the same strong BF attraction.
@@
Let us now analyze in detail the effect of the hybridization scale $\Delta_{0}$ through the energy spectrum $E_{\mathbf{P}}^{\pm}$ of the molecular quasi-particles defined by Eq. (29) of Sec. 3.1. In Fig. 4 the hybridized dispersions $E_{\mathbf{P}}^{\pm}$ are compared with those of the unhybridized molecular and atomic fermions for coupling $g=2.8$ and different values of the boson concentration $x$. At exactly matched densities ($x=1$) all fermions are essentially paired with all bosons ($n_{0}\approx 0$) and the system is effectively made of a gas of non-interacting molecules, with the unhybridized dispersions coinciding with the hybridized ones since $\Delta_{0}\simeq 0$.
@@
As soon as the concentration decreases (see the case $x=0.95$ in Fig. 4(a)), the hybridization energy scale $\Delta_{0}$ sharply rises and hybrid quasi-particles form out of the molecular and unpaired Fermi states. As the concentration further decreases, even though $n_{0}$ remains exponentially suppressed, the presence of a non-zero $\Delta_{0}\propto\sqrt{n_{0}{\varepsilon_{0}}}$ shifts down the energy dispersion $E_{\mathbf{P}}^{-}$ with respect to the unhybridized dispersion $\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}}$. This effect is rooted in the kinetic energy cost when filling up the paired and unpaired Fermi spheres: in order to reduce it, the molecular and atomic wave-functions hybridize quantum-mechanically by overlapping with the condensate wave-function (thus gaining in delocalization energy). The bottom panels of Fig. 4 display the typical level crossing (dotted and dashed lines) and avoided crossing (solid and dot-dashed lines) of a non-hybridized and hybridized two-level system, respectively. At the crossover concentration $x\simeq 0.7$, the hybridization is about maximal and the character of the occupied states (bottom pole dispersion) switches from molecular to mostly atomic as $x$ is reduced (with an effective mass approaching $m_{\mathrm{F}}$). At small concentration ($x=0.175$), the dispersion of occupied composite fermion states $E_{\mathbf{P}}^{-}$ approaches the non-interacting atomic dispersion $\xi_{\mathbf{P}}^{\mathrm{F}}$, albeit with a small quasi-particle weight $v_{\mathbf{P}}^{2}\simeq 0$ (see also Fig. 5). So, in this regime, the occupied composite fermion states are determined by the small participation of molecules to the hybridized dispersion $E_{\mathbf{P}}^{-}$ with a predominant atomic fermion character.
@@
Figure 5: Composite-fermion momentum distribution $n_{\mathrm{CF}}(\mathrm{P})=v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ as a function of $P/k_{\rm F}$, for BF coupling $g=2.8$, BB repulsion $\eta_{\mathrm{B}}=0$, and concentrations matching those of Fig. 4, along with the case $x=1$.